1. I am thinking of a number.If you add 10 to it and divide the sum by 5, you get 7. What number am i thinking of?

2. Divide my number by 5 and add 10 to the quotient.You get 15, what is my number?
3. If I divide my number of marbles among Miles, Mel and Rosie, each of them get 30 marbles.How many marbles do I have?
4. In the morning 15 cars parked in the basement of the building,At noontime,the number of cars
was doubled and then, at 4:00pm, 7 more cars wee parked.At 6:00pm,, 8 cars went out.How many cars were then still parked?
5. Add 12 to the number that I am thinking of. Then divide the sum by 3 and then, multiply the quotient by 5.Subtract 15 from the product you get 20.What number am I thinking of?
6. Milo, Ken and Vernon are car washers. Milo washed one-half of all the cars
that time. Ken washed one-forth of he remaining cars, and Vernon washed the rest. If Vernon washed 15 cars, how many cars did the three people wash?
7. Helen went to a store and spent half of her money plus 15 pesos more. She went to a second store and spent half of her remaining money plus 15 pesos more. Then she had exactly 75 pesos left. How much did she have at the beginning>
8. Mr. Ablir operates the elevator in a school building. He took Mr. Espe up 6 floors from the floor on which Mr Espe had his class. Then Mr. Ablir went down 9 floors to the first floor lobby. What is the number of the floor on which Mr. Espe had his class?

Answer by ReadingBoosters(3246)

1.) Work backward. 7 * 5 = 35 - 10 = 25 is the number
Proof:25 + 10 = 35/5 = 7
2.) Work backward. 15-10 = 5*5 = 25
Proof:25 / 5 = 5 + 10 = 15
3.) x/3 = 30 so, multiply both sides by 3 and x = 90
4.) 15*2 = 30+7=37-8=29
5.) Work backward. 20+15=35/5=7*3=21-12=9
Proof: 9+12=21/3=7*5=35-15=20
6.) x is the number of cars: (x - 1/2x)-1/4(x - 1/2x) = 15
x - 1/2x - 1/4x + 1/8x = 15
8/8x - 4/8x - 2/8x + 1/8x = 15
3/8x = 15
x = 40 cars
7.) 1/2(1/2x - 15) - 15 = 75
1/4x - 15/2 - 15 = 75
1/4x - 15/2 - 30/2 = 75
1/4x - 45/2 = 75
1/4x = 75 + 45/2
1/4x = 150/2 + 45/2
1/4x = 195/2
x = 195/2(4)
x=$390 starting with
Proof = 390(1/2)=195-15=180(1/2)90-15=75
Proof: (1/2)40=20 by Milo. 40-20=20,(1/4)20=5 by Ken 20-5=15 by Vernon
8.)x+6-9=0 so x-3=0 so x=3rd floor
Proof: 3 + 6 = 9 - 9 = 0